OpenStudy (is3535):

Andrew estimated the weight of his dog to be 60 lb. The dog’s actual weight was 68 lb. What was the percent error in Andrew’s estimate? Round your answer to the nearest tenth of a percent.

OpenStudy (is3535):

i think 12 or 11.7

OpenStudy (is3535):

@sammixboo

OpenStudy (is3535):

@Vocaloid

OpenStudy (is3535):

@butterflydreamer

OpenStudy (is3535):

@ro561man

OpenStudy (rajat97):

is it 11.76?

OpenStudy (is3535):

i think 12 or 11.7

OpenStudy (butterflydreamer):

11.76 is correct :) If we were to round it, it would become 11.8 :)

OpenStudy (dan815):

wrong

OpenStudy (is3535):

so 11.8 is my answerr

OpenStudy (anonymous):

divide it bye .10

OpenStudy (rajat97):

the qualified helper says its wrong

OpenStudy (butterflydreamer):

yes, that's correct

OpenStudy (is3535):

ther not smart the qualified helper

OpenStudy (rajat97):

i'll be back in some time

OpenStudy (anonymous):

i get 680

OpenStudy (dan815):

BRUH

OpenStudy (butterflydreamer):

LOL... um. I think what you need to do is first find the difference (error) between your estimated weight and the actual weight. So 60 - 68 = -8 Ignore the negative... so our error is 8. Then to find the percentage error we do : error/ actual value = 8 / 68 = 2/17 Then change it into a percentage by multiplying by 100. 2/17 * 100 = 11.76 Then round it to the nearest tenth ( 1 decimal point) 11.76 = 11. 8 (1dp)

OpenStudy (anonymous):

oh that makes since lol

OpenStudy (butterflydreamer):

.......................................................... http://prntscr.com/7bd9ro

OpenStudy (butterflydreamer):

sooooooooooo 11.8% is correct if you follow the steps above/ use the formula above :)